OpenStudy (anonymous):

What is the common ratio for the following geometric sequence? 9, 18, 36, 72, 144 does that mean greatest commmon factor

OpenStudy (heisenberg):

no a geometric sequence is a sequence where each term is the previous term times a common ratio. so: 18 = 9x 36 = 18x 72 = 36x ... and so on what is x?

OpenStudy (anonymous):

i do not get it

OpenStudy (heisenberg):

do you know what a sequence is?

OpenStudy (anonymous):

yes but how do i know wht the common ratio is?

OpenStudy (heisenberg):

choose a term, look at the previous term, and determine what needs to be multiplied with the first term to achieve the second term. in my equations above, you can solve for x and get the answer.

OpenStudy (anonymous):

23 is common raito ????????

OpenStudy (anonymous):

\[r = \frac{a_{n+1}}{a_n}\]

OpenStudy (anonymous):

is the ratio 1:2? Because every term is twice the one before it?

OpenStudy (anonymous):

how do i go about using that formula with what i am givin

OpenStudy (anonymous):

\[r = \frac{a_{n+1}}{a_n} = \frac{a_2}{a_1} = \frac{18}{9} = 2\]

OpenStudy (anonymous):

i though you do + 1 to a2 on top

OpenStudy (anonymous):

I think you should first understand what a sequence is.

OpenStudy (anonymous):

i do

OpenStudy (anonymous):

The numbers below a are what are known as subscripts.

OpenStudy (anonymous):

do u not add 1 to top?

OpenStudy (anonymous):

They represent the index of the term in the sequence.

OpenStudy (anonymous):

Geometric sequences have a common ratio which holds for any two given terms. The formula for the ratio expresses this fact.

OpenStudy (heisenberg):

\(a_1 =\) the 1st term of the sequence \(a_2 =\) the 2nd term of the sequence \(a_3 =\) the 3rd term of the sequence \(a_n =\) the nth term of the sequence \(a_{n+1} =\) the (n + 1)th term of the sequence so for any nth term, the nth plus 1 term is a multiple by what factor?

OpenStudy (anonymous):

help me with my other question